Find the charge in coulomb of 1g ion of N^3- .plz frwrd the solu as soon as possible

Grade:11

2 Answers

varun
6 years ago
HELLO APURV,
Charge on 1 atom of N^3- = 3e-
and
Mole of N^3- = 1/14
So moles of charge in 1gm of N^3- = 1
                                                              ----  X 3
                                                               14
So total charge on N^3- = 3  
                                           ----   X avogadro no
                                            14
NOW
1Na  ---------------> 96500C
3/14 Na --------------> 96500 X 3
                                               ---
                                              14
                                  => 20678.5
THANK YOU
Vaidhehi
5 years ago
One ion carries 3×1.6×10^19- coulombs 
Then 1 g ion N^3-(1 Mole)carries
                 =3×1.6×10^19-×6.2×10^23
                 =2.89×10^5couloms